Given an RC series circuit that has an emf source of 110 volts, a resistance of 3 kilo-ohms, a capacitance of 5 microfarad and the initial charge of the capacitor is 50 micro coulomb. What is the charge in the circuit at the end of 0.05 seconds?

1
Expert's answer
2021-12-20T15:35:53-0500

Differential equation of a LR circuit 


"Rq'+\dfrac{q}{C}=U"

Given "U=110V, R=3000\ Ohm, C=5\times10^{-6}\ F"

Substitute


"3000q'+200000q=110""q'+\dfrac{200}{3}q=0.11"

Integrating factor


"\mu(t)=e^{\int(200/3)dt}=e^{(200/3)t}"

"e^{(200/3)t}(q'+\dfrac{200}{3}q)=0.11e^{(200/3)t}"

"d(e^{(200\/3)t}q)=0.11e^{(200\/3)t}dt"

Integrate


"\int d(e^{(200/3)t}q) =\int 0.11e^{(200/3)t}dt"

"e^{(200\/3)t}q=0.00165e^{(200\/3)t}+C"

"q(t)=1.65\times 10^{-3}+Ce^{-(200/3)t}"

"q(0)=50\times 10^{-6} A"

"0.05\times 10^{-3} =1.65\times 10^{-3}+C"

"C=-1.60\times 10^{-3}"

"q(t)=1.65\times 10^{-3}-1.60\times 10^{-3}e^{-(200/3)t}"

"q(0.05)=1.65\times 10^{-3}-1.60\times 10^{-3}e^{-(200/3)(0.05)}"

"q(0.05)=1.593\ mC"